If p is the avg (arthemetic mean) of 3a and 7, q is the avg of 5a and 13, and r is the avg of 4a and 22, what is the average of p,q and r in terms of a?
3a+7
4a+10.5
2a+7
12a+42

Found 2 solutions by rothauserc, Theo:
Answer by rothauserc(4718)

(3a+7)/2 = p
:
(5a+13)/2 = q
:
(4a+22)/2 = r
:
(p+q+r)/3 = (((3a+7)/2) +((5a+13)/2) +((4a+22)/2))/3 =
:
((3a+7) +(5a+13) +(4a+22))/6 =
:
(12a+42)/6 =
:
2a+7
:
******************************
average of p, q and r is 2a+7
******************************
:

Answer by Theo(13342)

i think 2a + 7 is your answer.

p is the average of 3a + 7 means p = (3a + 7)/2

q is the average of 5a and 13 means q = (5a + 13)/2

r is the average of 4a and 22 means r = (4a + 22)/2

to get the average of p, q, and r, you need to add them together and divide by 3.

adding them all together, you get (12a + 42)/2 = 6a + 21.

divide 6a + 21 by 3 and you get 2a + 7.

that should be your answer.